# Low-Redundant Optimization for Large Language Model Alignment

### v1

@@ 1 Introduction
However, it is hard to train a well-aligned LLM for complex tasks and scenarios Feng et al. (2024); Gekhman et al. (2024). The key issue is that LLMs might overfit into the unexpected patterns or superficial styles in the human preference data (i.e., pairs of positive-negative response) Du et al. (2024). It is the side effect of their powerful learning capability derived from the large-scale trainable parameters Song et al. (2024); Meng et al. (2024a). Recently, a surge of work Frankle and Carbin (2019); Wang et al. (2024b) has found that each neuron (i.e., one of the trainable values of the parameter matrixes in LLMs) is relevant with special knowledge, and the neurons in LLMs are generally sparsely activated. Inspired by it, we consider whether the irrelevant neurons exist in LLMs during the alignment process and the full-parameter trained LLMs might lead to redundant updates on alignment-irrelevant neurons. Thus, we conduct the empirical experiment using DPO algorithm, where we only update the top/last-10% neurons according to their accumulated gradient values, which indicate the relevance between the current neuron and the downstream scenario Pruthi et al. (2020). As shown in Figure 1, with the top-10% trainable neurons, LLMs can converge faster and achieve better performance than optimizing all the neurons. It indicates the existence of alignment-irrelevant neurons and redundant updates in DPO training, affecting the convergence and final performance. Besides the redundant neurons, previous work Lin et al. (2024) has shown the existence of redundant tokens in training data.
@@ 2 Related Work
Large Language Models. LLMs have shown remarkable performance on various tasks qwe (2024); Meta (2024); Javaheripi et al. (2023). Generally, the training process of LLMs includes three stages, i.e., pre-training, supervised fine-tuning (SFT), and alignment Ouyang et al. (2022); Touvron et al. (2023). In the training process, previous work has selected valuable data to train the LLMs via leveraging gradient Xia et al. (2024) or perplexity Lin et al. (2024); Xie et al. (2023), Besides, synthetic training data from powerful LLMs (e.g., GPT-4, Claude 3) has been widely utilized for improving the weak LLMs Xu et al. (2023); Ben Allal et al. (2024); Liu et al. (2024), especially for specific scenarios (e.g., mathematical tasks or code synthesis tasks) Yue et al. (2023); Zhou et al. (2024). However, given the large expenses of the LLM training, existing work Hu et al. (2022); Li and Liang (2021); Dettmers et al. (2023) has revealed that training only a small number of the parameters can achieve comparable performance with whole-parameters training. In this work, we focus on the alignment stage and leverage the low-redundant optimization to improve the existing LLMs.

### v2

# Not Everything is All You Need: Toward Low-Redundant Optimization for Large Language Model Alignment

@@ 1 Introduction
However, it is hard to train a well-aligned LLM for complex tasks and scenarios Feng et al. (2024); Gekhman et al. (2024). The key issue is that LLMs might overfit into the unexpected patterns or superficial styles in the human preference data (i.e., pairs of positive-negative response) Du et al. (2024a). It is the side effect of their powerful learning capability derived from the large-scale trainable parameters Song et al. (2024); Meng et al. (2024a). Recently, a surge of work Frankle and Carbin (2019); Wang et al. (2024b) has found that each neuron (i.e., one of the trainable values of the parameter matrixes in LLMs) is relevant with special knowledge, and the neurons in LLMs are generally sparsely activated. Inspired by it, we consider whether the irrelevant neurons exist in LLMs during the alignment process and the full-parameter trained LLMs might lead to redundant updates on alignment-irrelevant neurons. Thus, we conduct the empirical experiment using DPO algorithm, where we only update the top/last-10% neurons according to their accumulated gradient values, which indicate the relevance between the current neuron and the downstream scenario Pruthi et al. (2020). As shown in Figure 1, with the top-10% trainable neurons, LLMs can converge faster and achieve better performance than optimizing all the neurons. It indicates the existence of alignment-irrelevant neurons and redundant updates in DPO training, affecting the convergence and final performance. Besides the redundant neurons, previous work Lin et al. (2024) has shown the existence of redundant tokens in training data.
@@ 2 Related Work
Large Language Models. LLMs have shown remarkable performance on various tasks qwe (2024); Meta (2024); Javaheripi et al. (2023). Generally, the training process of LLMs includes three stages, i.e., pre-training, supervised fine-tuning (SFT), and alignment Ouyang et al. (2022); Touvron et al. (2023). In the training process, previous work has selected valuable data to train the LLMs via leveraging gradient Xia et al. (2024) or perplexity Lin et al. (2024); Xie et al. (2023), Besides, synthetic training data from powerful LLMs (e.g., GPT-4, Claude 3) has been widely utilized for improving the weak LLMs Xu et al. (2023); Ben Allal et al. (2024); Liu et al. (2024), especially for specific scenarios (e.g., mathematical tasks or code synthesis tasks) Yue et al. (2023); Zhou et al. (2024b). However, given the large expenses of the LLM training, existing work Hu et al. (2022); Li and Liang (2021); Dettmers et al. (2023) has revealed that training only a small number of the parameters can achieve comparable performance with whole-parameters training. In this work, we focus on the alignment stage and leverage the low-redundant optimization to improve the existing LLMs.
@@
Low-Cost LLM Training. Due to the high computation cost of LLM training, previous work either trains a small portion of parameters in LLMs, or trains small-scale LLMs. For the former, existing work has shown the parameter redundancy in LLMs, and proposed related approaches to reduce the trainable parameters in LLM Hu et al. (2022); Li and Liang (2021). Besides, quantization has also been used to further reduce the requirements of GPU memory Dettmers et al. (2023). Moreover, recent work has further improved the above methods by adjusting the updated parameters within LLMs or improving the training strategies Zhou et al. (2024a); Du et al. (2024b). For training small LLMs, to break the ceiling derived from the scaling law, knowledge distillation methods have been widely used for transferring the capabilities from the teacher model to small LLMs He et al. (2023). However, it is also promising to small models for hyper-parameters searching, to obtain the optimal training hyper-parameters for large LLMs Hu et al. (2024). In this work, we mainly focus on training a small portion of parameters to reduce the useless redundancy in the human alignment training process. Such a way is capable of reducing the training cost and also improves the performance.
@@ 6 Conclusion
Acknowledgement
This work was partially supported by National Natural Science Foundation of China under Grant No. 62222215, Beijing Natural Science Foundation under Grant No. L233008 and 4222027. Xin Zhao is the corresponding author.